工業英語 2 単位 (選択)

Communication using Technical English  コインカー, パンカジ · マドゥカー · 特任助教 / 先端技術科学教育部

【授業目的】 To improve students’ ability to write and speak on technical subjects

using English. 【授業概要】 This is a language course designed for engineering students. It’s aim

is to increase students’ confidence communicating in English. The focus will be on practicing simple technical language that is commonly used in science and engineering. Vocabulary and communication techniques are introduced and explored through the use of real-life science topics. This is a practical course, which requires a high degree of student participation.

【キーワード】リーディング，ライティング，スピーキング，リスニング【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】 None 【履修上の注意】 None 【到達目標】 The goal of this course is to increase the students’ technical English

vocabulary and confidence communicating in English. By the end of the course students should be able to give an oral presentation and written summary of a technical topic in English.

【授業計画】 1. Course outline and self-introductions

2. Language for spatial description

3. Basic numbers in science

4. Body language and presentation practice

5. Scientific units of measurement

6. Description and cause-and-effect

7. Compare and contrast

8. Presentation techniques, and practice

9. More presentation practice

10. Definition and description

11. Structure, organization, explanation

12. Visual aids and science

13. Final presentations: assessment

14. Final presentations: assessment

15. Final presentations: assessment 【成績評価基準】 Grades will be determined by two assessments: an oral presentation

(40%) and a written essay (60%) on a technical topic. 【教科書】 Technical English 1, David Bonamy, Pearson-Longman 【参考書】 Presenting Science (1st and 2nd editions), MacMillan Languagehouse.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215855 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ コインカー (イノベーションセンター 3 階, 088-656-7643, koinkar@cicee.t

okushima-u.ac.jp) ���� 【備考】

⋄ An English-Japanese dictionary is also recommended.

⋄ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

Communication using Technical English 2 units (selection)

Pankaj Koinkar · Designated Assistant Professor / Graduate School of Advanced Technology and Science

Target⟩ To improve students’ ability to write and speak on technical subjects using

English. Outline⟩ This is a language course designed for engineering students. It’s aim

is to increase students’ confidence communicating in English. The focus will be on practicing simple technical language that is commonly used in science and engineering. Vocabulary and communication techniques are introduced and explored through the use of real-life science topics. This is a practical course, which requires a high degree of student participation.

�������⟩ reading, writing, speaking, listening Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ None Notice⟩ None Goal⟩ The goal of this course is to increase the students’ technical English

vocabulary and confidence communicating in English. By the end of the course students should be able to give an oral presentation and written summary of a technical topic in English.

Schedule⟩

1. Course outline and self-introductions

2. Language for spatial description

3. Basic numbers in science

4. Body language and presentation practice

5. Scientific units of measurement

6. Description and cause-and-effect

7. Compare and contrast

8. Presentation techniques, and practice

9. More presentation practice

10. Definition and description

11. Structure, organization, explanation

12. Visual aids and science

13. Final presentations: assessment

14. Final presentations: assessment

15. Final presentations: assessment Evaluation Criteria⟩ Grades will be determined by two assessments: an oral

presentation (40%) and a written essay (60%) on a technical topic. Textbook⟩ Technical English 1, David Bonamy, Pearson-Longman Reference⟩ Presenting Science (1st and 2nd editions), MacMillan Languagehouse.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215855

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Koinkar (イノベーションセンター 3 階, +81-88-656-7643, koinkar@cicee.t

okushima-u.ac.jp) ���� Note⟩

⋄ An English-Japanese dictionary is also recommended.

⋄ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．